OpenStudy (anonymous):

Under ideal conditions a certain bacteria population is known to double every three hours. Suppose that there are initially 70 bacteria. (a) What is the size of the population after 9 hours? bacteria (b) What is the size of the population after t hours? bacteria (c) What is the size of the population after 16 hours? (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

a) Double every 3 hours -> 70 *2*2*2 (*2 every 3 hours so 3 times)

OpenStudy (anonymous):

Now generalize it for t and use your formula to do part c

OpenStudy (anonymous):

okaay thanks you

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

\[ \text{We can write the population as a function of time as follows.}\\ \ \ \ P(t)=P_02^\frac{t}3\\ \text{where}\\ \ \ \ P(t)\text{ - the population after }t\text{ hours}\\ \ \ \ \ P_0\ \text{ - the initial population}\\ \text{We're given as follows.}\\ \ \ \ P_0=70\\ \text{So we can reduce our function to...}\\ \ \ \ P(t)=70\times2^\frac{t}3\\ \text{a. }P(9)=70\times2^\frac93=70\times2^3=560\\ \text{b. }P(t)=70\times2^\frac{t}3\\ \text{c. }P(16)=70\times2^\frac{16}3\simeq2822 \]

OpenStudy (anonymous):

thanks you every much

OpenStudy (anonymous):

Estimate the time for the population to reach 20,000. (Round your answer to two decimal places.)

OpenStudy (anonymous):

ask oldrin.bataku she will give u answer.....

OpenStudy (anonymous):

\[ \text{We need to find }t\text{ such that }P(t)\text{ reaches 20,000.}\\ \ \ \ P(t)=20000\\ \text{Let's try to reduce for }t.\\ \ \ \ 70\times2^\frac{t}3=20000\\ \ \ \ 2^\frac{t}3=\frac{2000}7\\ \ \ \ \frac{t}3=\log_2\frac{2000}7\\ \ \ \ t=3\ \log_2\frac{2000}7\\ \ \ \ \ \simeq24.48\\ \text{Therefore it should take roughly 24.48 hours.} \]

OpenStudy (anonymous):

PS @estudier I'm a boy.

OpenStudy (anonymous):

@oldrin.bataku OK, I believe u.....